OpenStudy (anonymous):

find the average rate of change for the following function on the indicated interval. f(x)=3sqrt(x) ; (4, 25)

OpenStudy (anonymous):

average is \[\frac{1}{25}\int_4^{25} 3\sqrt{x}dx\]

OpenStudy (anonymous):

that is a mistake

OpenStudy (anonymous):

\[\frac{1}{21}\int_4^{25} 3\sqrt{x}dx\]

OpenStudy (anonymous):

length of path from 4 to 25 is 21, not 25

OpenStudy (anonymous):

might as well write \[\frac{1}{7}\int_4^{25}\sqrt{x}dx\] anti derivative is easy enough, it is \[\frac{2}{3}\sqrt{x^3}\] plug in 25, 4, and subtract. ok from there or you need it worked out?

OpenStudy (anonymous):

i think i'm good from there. i just plug 25 and 4 in for x, correct?

OpenStudy (anonymous):

right. and don't forget the \[\frac{1}{7}\] out front. it will be \[\frac{1}{7}\times \frac{2}{3} (\sqrt{25}^3-(\sqrt{4})^3)\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

quick question: would i do the same thing for f(x)=x^3-2x ; (0,4)